MEGHÍVÓ

2003. március 17-én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

Jerry L. Kazdan (University of Pennsylvania/CEU)

SOME NONLINEAR PROBLEMS IN GEOMETRY

című előadása.

One seemingly elementary question in differential geometry asks if is one can always embed a smooth two-dimensional Riemannian manifold in R^{3} – at least locally. This and other related problems lead to nonlinear partial differential equations (”Monge-Ampère equations”) that are at the cutting edge where classical techniques don’t apply and we have little intuition.

In this lecture I’ll describe some recent significant advances. I will not assume technical understanding of either partial differential equations or differential geometry.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!

2003. március 17-én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

Jerry L. Kazdan (Pennsylvania State University/CEU)

SOME NONLINEAR PROBLEMS IN GEOMETRY

című előadása.

One seemingly elementary question in differential geometry asks if is one can always embed a smooth two-dimensional Riemannian manifold in R^{3} – at least locally. This and other related problems lead to nonlinear partial differential equations (”Monge-Ampère equations”) that are at the cutting edge where classical techniques don’t apply and we have little intuition.

In this lecture I’ll describe some recent significant advances. I will not assume technical understanding of either partial differential equations or differential geometry.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!